# ID-Composer: Multi-Subject Video Synthesis with Hierarchical Identity Preservation

### v2

@@ 5 Conclusion
In this work, we introduced ID-Composer, an innovative framework for multi-subject video generation that integrates hierarchical identity-preserving attention and a pretrained Vision-Language Model into a latent diffusion backbone. Moving beyond simple feature concatenation, ID-Composer captures complex intra- and inter-subject interactions, preserves identities, incorporates semantics, and ensures temporal consistency. Extensive experiments demonstrate that ID-Composer achieves state-of-the-art performance, outperforming existing methods across a comprehensive suite of metrics. Our work marks an important step towards general and scalable multi-subject video generation in complex, real-world scenarios.
Limitations and Future Work
While ID-Composer excels at identity preservation and multi-subject video generation, it remains limited by high computational costs and difficulties with complex interactions and fine-grained dynamics. Future work includes developing more efficient architectures, incorporating physics-aware priors and 3D reasoning, mitigating biases in pretrained components, and advancing controllable generation for fine-grained attribute, action, and interaction specification.
Appendix A Declaration of LLM Usage
During the writing of the manuscript, we utilized a Large Language Model (ChatGPT) as a writing assistant. The scope of its usage was limited to improving grammar, polishing sentences, and enhancing the clarity and fluency of this manuscript. The method, claims, experimental results and conclusions are developed by the authors.
Appendix B More Implementation Details
Reproducibility
To facilitate reproducibility, we present our detailed experimental settings and evaluation metrics in Section 4.1. This section provides a comprehensive description of our implementation details. Moreover, our source code and pre-trained models will be publicly available.
B.1 Details of Network
We use a pre-trained VAE model from Wan-Video (Wan, 2025) to convert images from pixel space to latent space and vice versa. The latent representation has a down-sample ration of 8 and a latent channel of 16, and is then processed by a 2 × 2 patch embedding layer to reduce the spatial size and match the hidden dimension of the LLM backbone. The VAE model is frozen during training.
Network Architecture
Our video generation model is built upon three core modules: a Vision-Language Model (VLM) (Bai et al., 2025) for multi-modal comprehension, a Variational Autoencoder (VAE) for spatial compression, and a Multi-Modal Diffusion Transformer (DiT) (Wan, 2025) serving as the latent video diffusion backbone. The VLM, with a total of 7B parameters, is responsible for encoding textual prompts and subject images into a shared semantic space. It consists of a 32-layer Vision Transformer (ViT) and a 28-layer Large Language Model (LLM). The VAE is designed to encode video frames into a compact latent representation and subsequently decode them back to the pixel space. Its encoder and decoder utilize a spatial scale factor of $8\times 8$, managing a parameter count of 54M and 73M, respectively. The core of our generative process is the DiT, a substantial 14B parameter transformer that operates on the latent codes produced by the VAE. With 60 layers and an intermediate size of 12,288, it is architected to model complex spatio-temporal dynamics for high-fidelity video synthesis. A detailed breakdown of the architectural configurations for each component is provided in Table 3.
Table 3: Detailed configuration of our model’s primary architectural components. The VLM, VAE, and MMDiT modules are designed to handle multimodal understanding, spatial compression, and latent diffusion, respectively.
Configuration VLM VAE MMDiT
@@ B.1 Details of Network
# Layers 32 28 11 15 60
# Num Heads (Q / KV) 16 / 16 28 / 4 - - 24 / 24
Head Size 80 128 - - 128
Intermediate Size 3,456 18,944 - - 12,288
Patch / Scale Factor 14 - $\times8$ 8 $\times$ 8 2
@@
# Parameters 7B 54M 73M 1.3B / 14B
B.2 Details of Reward Calculation
Appendix C Dataset Curation Details
C.1 Dataset Statistics
@@
Figure 7 illustrates the composition and distribution of scenarios in our training dataset, which is meticulously curated from three primary sources to ensure comprehensive coverage and high fidelity. The foundational component is the large-scale OpenS2V dataset (Yuan et al., 2025), which contributes a substantial volume of 218,230 videos, accompanied by 535,259 masked subject images and 3,098 generated counterparts. To further elevate the dataset’s quality, we incorporate a high-quality collection, comprising 9,374 pristine videos and an additional 3,000 generated subject images. Finally, to provide rich priors for appearance synthesis, we integrate a dedicated Subject Image dataset. This static collection contains 14,976 images, which are further categorized into 2,877 human-centric images and 3,458 isolated clothing items.
Figure 7: Statistics of the constructed dataset. The dataset is organized into four primary scenarios: Human, Objects, Environment, and Nature, each containing a variety of subcategories.
Figure 8: Statistics of the constructed dataset. (a) A word cloud illustrates the rich linguistic diversity of the reference images. The video duration distributions are shown for the two sources: (b) OpenS2V and (c) our curated videos.
C.2 Comparison of Existing Datasets
@@
We analyze various strategies for constructing the reference datasets. As shown in Figure 9, Naive data augmentation offers limited visual diversity and is prone to generating occluded subjects, which compromises generalization to complex scenes. Alternatively, employing on-the-shelf image-consistent generators, such as Flux-Kontext (Labs et al., 2025) or even GPT-4o (Hurst et al., 2024), often introduces undesirable appearance artifacts and fails to maintain subject fidelity.
In contrast, our approach, which leverages a tailored in-context learning prompt for the Gemini-Flash-Image model (Fortin et al., 2025), excels at subject decomposition and synthesis, yielding more faithful and varied reference images.
Figure 9: Comparison of reference image construction strategies.
Appendix D Details of Evaluation Metrics
@@
NexusScore: This metric quantifies the subject consistency between a generated video $\mathbf{V}=\{f_{1},\dots,f_{T}\}$ and a reference image $\mathbf{I}_{\text{ref}}$. It employs a two-stage pipeline that first leverages Grounded-SAM to generate a subject mask $M_{t}$ for each frame $f_{t}$. Subsequently, a refined CLIP-based image encoder $\mathcal{E}_{img}$ is used to compute the cosine similarity between the feature embedding of the reference image and that of the masked subject in each frame. This targeted approach ensures the evaluation focuses strictly on identity fidelity by mitigating background interference. The final score is the average similarity over all $\mathbf{T}$ frames, formulated as: $S_{\text{Nexus}}=\frac{1}{\mathbf{T}}\sum_{t=1}^{\mathbf{T}}\frac{\mathcal{E}_{\text{img}}(\mathbf{I}_{\text{ref}})\cdot\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})}{\|\mathcal{E}_{img}(\mathbf{I}_{\text{ref}})\|\|\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})\|}$ where $\odot$ denotes element-wise multiplication.
@@
NaturalScore: This metric assesses the perceptual realism and physical plausibility of a generated video $\mathbf{V}$. We utilize a state-of-the-art Multimodal Large Language Model (MLLM) to perform a deep semantic analysis. The model is presented with the video and a carefully designed prompt which instructs it to evaluate spatio-temporal consistency, identify violations of physical laws, and detect visual artifacts characteristic of generative models. The model’s direct output, a normalized score, serves as the final measure of the video’s naturalness.
@@
GmeScore: This metric evaluates the semantic alignment between a generated video $\mathbf{V}$ and its corresponding text prompt $\mathbf{C}$, proposed as an enhanced alternative to conventional CLIPScore. To overcome the limitations of prior methods in handling long-form text, GmeScore leverages the advanced text comprehension capabilities of the Qwen2-VL MLLM (Bai et al., 2025), which provides a powerful text encoder $\mathcal{E}_{text}$ and a temporally-aware video encoder $\mathcal{E}_{video}$. Unlike methods that average per-frame similarities, GmeScore computes a global feature representation for the entire video, thereby capturing its holistic temporal dynamics. The final score is the cosine similarity between the global video embedding and the text embedding: $S_{\text{Gme}}=\frac{\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\cdot\mathcal{E}_{\text{text}}(C)}{\|\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\|\|\mathcal{E}_{\text{text}}(C)\|}$ This formulation provides a more accurate and comprehensive assessment of text-to-video relevance, especially for complex narratives.
Appendix E More Comparison Results

### v3

@@ 5 Conclusion
In this work, we introduced ID-Crafter, an innovative framework for multi-subject video generation that integrates hierarchical identity-preserving attention and a VLM into a latent diffusion backbone. Moving beyond simple feature concatenation, ID-Crafter captures complex intra- and inter-subject interactions, preserves identities, incorporates semantics, and ensures temporal consistency. Extensive experiments demonstrate that ID-Crafter achieves state-of-the-art performance, outperforming existing methods across comprehensive metrics. Our work marks an important step towards general and scalable multi-subject video generation in complex, real-world scenarios. Limitations and Future Work Despite its strength in multi-subject identity preservation, ID-Crafter has limitations in modeling complex interactions and fine-grained dynamics. Future work will focus on incorporating physics-aware priors, mitigating biases in pretrained components, and advancing fine-grained controllable generation of attributes, actions, and interactions.
[…]
Supplementary Material
Appendices Overview
The supplementary material includes the following sections:
  • Sec. A: More Implementation Details.
    • Details of Network Architecture (Sec. A.1).
    • Details of Hierarchical Identity-Preserving Attention (Sec. A.1).
    • Details of Reward Calculation (Sec. A.2).
  • Sec. B: Dataset Curation Details.
    • Dataset Composition and Statistics (Sec. B.1).
    • Reference Image Construction Strategies (Sec. B.2).
  • Sec. C: Evaluation Metric Details.
  • Sec. D: Additional Results and Analysis.
    • Evaluation on Multi-Subject Scenarios (Sec. D.1).
    • Human Preference Study (Sec. D.2).
  • Sec. E: Additional Discussion.
    • Computational Overhead (Sec. E.1).
    • Ethics Statement (Sec. E.2).
Appendix A More Implementation Details
Table S.4: Detailed configuration of our model’s primary architectural components. The VLM, VAE, and DiT modules are designed to handle multimodal understanding, spatial compression, and latent diffusion, respectively.
Configuration VLM VAE DiT
@@ Appendix A More Implementation Details
# Layers 32 28 11 15 30/40
# Num Heads (Q / KV) 16 / 16 28 / 4 - - 12 / 40
Head Size 80 128 - - 128/128
Intermediate Size 3,456 18,944 - - 8,960/13,824
Patch / Scale Factor 14 - $\times8$ $\times8$ -
@@
[…]
Table S.5: Ablation of Hierarchical Attention. We analyze the contribution of each stage in our hierarchical attention mechanism. The full three-stage model provides the best performance, with Stage 3 (cross-modal attention to VLM) being the most critical component for semantic alignment.
Method FaceSim $\uparrow$ NexusScore $\uparrow$ Total Score $\uparrow$
[…]
w/o Stage 1 (Intra-Subject) 63.2% 44.2% 54.1%
w/o Stage 2 (Inter-Subject) 64.5% 44.5% 54.5%
w/o Stage 3 (Cross-Modal) 58.9% 38.7% 50.8%
Vanilla Attention (Stage 3 only) 55.1% 39.1% 49.7%
A.1 Network Architecture
Our model’s architecture is founded on three synergistic modules: a Vision-Language Model (VLM), specifically the Qwen2.5-VL 7B variant [Qwen2.5-VL], for advanced multi-modal comprehension; a Variational Autoencoder (VAE) [rombach2022high] for efficient spatial compression; and a Diffusion Transformer (DiT) [peebles2023scalable, wan2025], which serves as the latent video diffusion backbone. The architectural configurations for each component are detailed in Table S.4.
The 7B-parameter VLM, comprising a 32-layer Vision Transformer (ViT) and a 28-layer Large Language Model (LLM), is responsible for encoding textual prompts and subject images into a unified, semantically rich space. The VAE, with a 54M-parameter encoder and a 73M-parameter decoder, utilizes an $8\times$ spatial downsampling factor, compressing video frames into a compact latent representation that is computationally tractable.
Subject Image Encoder
We employ the pre-trained VAE from Wan-Video [wan2025] to transform subject images from the pixel space into the latent space. This process yields a $16$-channel latent representation with an $8\times$ downsampling factor. This latent representation is subsequently processed by a $2\times 2$ patch embedding layer, which further reduces its spatial dimensions to align with the hidden dimension of the DiT backbone, ensuring seamless integration between the visual and generative components. The VAE weights are kept frozen during training to preserve its high-fidelity reconstruction capabilities.
Hierarchical Identity-Preserving Attention
As detailed in the main paper, our Hierarchical Identity-Preserving Attention module is central to preventing identity leakage in multi-subject generation. It comprises three sequential stages applied within each block of the DiT, designed to progressively refine and integrate identity features. Let $Z_{l}$ be the video latent tokens at layer $l$, and $F_{subj}=\{f_{1},f_{2},\dots,f_{N}\}$ be the token sets for the $N$ subject images.
Stage 1: Intra-Subject Attention.
To refine the feature representation for each subject independently, we first apply a standard self-attention layer [vaswani2017attention] to each subject’s token set $f_{i}$. This initial stage allows the model to capture fine-grained, identity-specific details and consolidate features for each subject before any cross-subject interaction occurs.
$f^{\prime}_{i}=\text{SelfAttention}(f_{i})\qquad\forall i\in\{1,\dots,N\}$ (S.9)
Stage 2: Gated Inter-Subject Attention.
To model interactions between subjects while strictly mitigating identity bleed, we employ a novel gated cross-attention mechanism. For each subject $f^{\prime}_{i}$, it queries the features of all other subjects $\{f^{\prime}_{j}\}_{j\neq i}$, which then modulates the output of the cross-attention.
$f^{\prime\prime}_{i}=f^{\prime}_{i}+\sigma(W_{g}f^{\prime}_{i})\odot\text{CrossAttention}(Q=f^{\prime}_{i},KV=\{f^{\prime}_{j}\}_{j\neq i})$ (S.10)
Stage 3: Cross-Modal Attention.
Finally, the refined subject tokens $F^{\prime\prime}_{subj}$ and the textual context tokens $f_{txt}$ attend to the main video latent tokens $Z_{l}$. This stage injects the rich semantic guidance from the VLM, now conditioned on the robust and disentangled subject identities. While the VLM features [Qwen2.5-VL] are computed only once, their influence is dynamic, as they modulate the video features at each layer of the DiT.
$Z^{\prime}_{l}=Z_{l}+\text{CrossAttention}(Q=Z_{l},KV=[F^{\prime\prime}_{subj};f_{txt}])$ (S.11)
Table S.5 provides a detailed ablation study on this design. These three stages form a single, cohesive block that is inserted within the first 20 layers of the DiT backbone. This design choice is based on the observation that identity features are most critical in the early stages of the generation process.
A.2 Details of GRPO-based Post-training
This section provides a more detailed breakdown of the reward calculation and hyperparameter settings for our online post-training phase, which leverages the Groupwise Policy Optimization (GRPO) algorithm [flowgrpo].
Advantage Normalization
To stabilize the training process and reduce variance in the policy gradient updates, we normalize the advantage term. The normalized advantage $\hat{A}_{i}$ for the $i$-th sample in a group of size $G$ is calculated from the total trajectory reward $r_{i}$:
$\hat{A}_{i}=\frac{r_{i}-\text{Mean}(\{r_{j}\}_{j=1}^{G})}{\text{Std}(\{r_{j}\}_{j=1}^{G})+10^{-8}}$ (S.12)
where the total reward $r_{i}$ is computed by averaging the frame-level rewards over the video’s duration to produce a single scalar value. The small epsilon ($10^{-8}$) is added for numerical stability. The constant $a$ for noise injection $\sigma_{t}$ was set to 1.0.
Reward Design
As defined in the main paper, the objective is to maximize a composite reward function $\mathcal{R}_{\text{total}}$, defined as a weighted sum of fidelity and quality rewards:
$\mathcal{R}_{\text{total}}(\mathbf{V}):=w_{\text{fid}}\mathcal{R}_{\text{fid}}(\mathbf{V},\mathcal{I})+w_{\text{qual}}\mathcal{R}_{\text{qual}}(\mathbf{V})$ (S.13)
We empirically set the weights to $w_{\text{fid}}=0.6$ and $w_{\text{qual}}=0.4$ to prioritize identity preservation while ensuring high quality. The fidelity reward $\mathcal{R}_{\text{fid}}$ is a combination of face-specific and holistic subject consistency scores, while the quality reward $\mathcal{R}_{\text{qual}}$ assesses aesthetic appeal and physical plausibility.
Hyperparameters and Sensitivity
The key hyperparameters used during the GRPO training are detailed in Table S.6. We also conducted a sensitivity analysis on critical parameters, with results presented in Table S.7, to demonstrate the robustness of our training setup.
Table S.6: Reinforcement Learning Hyperparameters. Key hyperparameters used for the online optimization stage with our GRPO algorithm.
Hyperparameter Value
Algorithm Flow-GRPO [flowgrpo, Deepseek-R1, shao2024deepseekmath]
Policy Learning Rate 1e-5
Optimizer AdamW
[…]
AdamW $\epsilon$ 1e-8
Batch Size 32
Max Gradient Norm 1.0
Training Dynamics and Reward Effectiveness
The effectiveness of our multi-faceted reward model is visualized in Figure S.7, which illustrates the evolution of key metrics during the online GRPO phase. The plot clearly shows a positive trend for identity fidelity (e.g., FaceSim), perceptual quality (Aesthetics), and Q-Align [wu2023qalign] over the course of training. This concurrent improvement across diverse axes validates that our reward model successfully guides the generator towards producing videos that are not only more identity-consistent but also more visually appealing and semantically accurate, without suffering from catastrophic forgetting or reward over-optimization on a single metric.
Ablation on Fidelity Reward
A crucial component of our total reward is the fidelity term $\mathcal{R}_{\text{fid}}$, which contains a balance factor $\gamma$ for the face-specific reward $\mathcal{R}_{\text{face}}$. Table S.7 presents an ablation study on the choice of this balance factor. The results indicate that a value of $0.5$ provides an optimal balance between average and worst-case identity fidelity, effectively preventing scenarios where one subject’s identity collapses while others are preserved.
Table S.7: Ablation on Fidelity Reward $\gamma$. We study the effect of $\gamma$ in $\mathcal{R}_{fid}$, which balances average and minimum identity similarity. A value of 0.5 provides the best trade-off.
$\gamma$ Avg FaceSim $\uparrow$ Min FaceSim $\uparrow$ Total Score $\uparrow$
0.0 (Average only) 67.2% 45.1% 54.8%
0.25 66.8% 50.3% 55.0%
0.5 66.1% 53.5% 55.2%
0.75 65.2% 52.1% 54.9%
1.0 (Minimum only) 63.9% 51.5% 54.5%
Figure S.7: Performance Curves of Online GRPO. We plot the moving average of key evaluation metrics over 800 training steps. Our method demonstrates simultaneous improvement in identity fidelity (FaceSim), perceptual quality (Aesthetics), and text-prompt alignment (e.g., Q-Align [wu2023qalign]), validating the effectiveness of our multi-faceted reward strategy.
Appendix B Dataset Curation Details
B.1 Dataset Composition and Statistics
@@
Our dataset is meticulously curated from three heterogeneous primary sources to ensure comprehensive coverage and high fidelity. The foundational component is the large-scale OpenS2V-Nexus dataset [yuan2025opens2v], which offers a wide array of authentic scenes and actions. To ensure the temporal coherence of the clips, we incorporate a video shot detection stage [pyscenedetect], which contributes a substantial volume of 218,230 videos, accompanied by 535,259 masked subject images and 3,098 generated counterparts. To further elevate the dataset’s quality, we incorporate a high-fidelity collection of professionally shot videos with detailed annotations, comprising 9,374 pristine videos and an additional 3,000 generated subject images. This synthetically generated data, where subjects rendered by cutting-edge image editing models are placed into novel contexts, systematically increases the diversity of subject-background compositions. Finally, to provide rich priors for appearance synthesis, we integrate a dedicated Subject Image dataset. This static collection contains 14,976 images, which are further categorized into 2,877 human-centric images and 3,458 isolated clothing items. The overall composition is summarized in Table S.8, and the entire data curation pipeline is illustrated in Figure S.8.
Table S.8: Detailed Composition of the Curated Training Dataset. This table provides a breakdown of the three primary data sources used to train ID-Crafter, detailing the content type, quantity, and its strategic role in the model’s training.
Dataset Source Content Type Quantity Primary Purpose
[…]
Generated Subjects 3,098 Initial subject synthesis capability
High-Quality Internal Collection High-Fidelity Videos 9,374 Fine-tuning for video quality
Generated Subjects 3,000 Enhancing appearance diversity
[…]
- Human-centric Images 2,877
- Isolated Clothing Items 3,458
Figure S.8: Overview of the Data Curation Pipeline. Our comprehensive data pipeline processes videos and images from diverse sources. It includes subject segmentation, high-quality reference image synthesis via an in-context learning prompt with advanced generative models, and rigorous filtering. This process yields a large-scale, high-fidelity dataset tailored for training identity-preserving video generation models.
B.2 Reference Image Construction Strategies
@@
We analyze various strategies for constructing reference datasets in Figure S.11. Naive data augmentation offers limited visual diversity and is prone to generating occluded subjects, which compromises generalization to complex scenes. Alternatively, employing on-the-shelf image-consistent generators, such as Flux-Kontext [labs2025flux] or GPT-4o [hurst2024gpt], often introduces undesirable appearance artifacts and, most critically, fails to maintain subject fidelity. In contrast, our approach, which leverages a tailored in-context learning prompt for the latest Gemini-Flash-Image model (a.k.a., Nano Banana) [nanobanana], excels at subject decomposition and synthesis, yielding more faithful and varied reference images.
Figure S.9: Statistics of the constructed dataset. The dataset is organized into four primary scenarios: Human, Objects, Environment, and Nature, each containing a variety of subcategories.
Figure S.10: Statistics of the constructed dataset. (a) A word cloud illustrates the rich linguistic diversity of the reference images. The video duration distributions are shown for the two sources: (b) OpenS2V and (c) our curated videos.
Figure S.11: Comparison of reference image construction strategies. In contrast to direct data augmentation or image editing models like Flux-Knotext [labs2025flux] and GPT-4O [hurst2024gpt], our data curation pipeline [Qwen2.5-VL, nanobanana] generates subject images with coherent multi-subject compositions.
Appendix C Evaluation Metric Details
Our evaluation protocol is principally derived from the OpenS2V-Nexus Benchmark [yuan2025opens2v]. We adopt their prescribed prompts but substitute the proprietary models with publicly available counterparts to facilitate evaluation and ensure reproducibility.
@@ Appendix C Evaluation Metric Details
NexusScore: This metric quantifies the subject consistency between a generated video $\mathbf{V}=\{f_{1},\dots,f_{T}\}$ and a reference image $\mathbf{I}_{\text{ref}}$. It employs a two-stage pipeline: (1) Grounded-SAM [ren2024grounded] generates a subject mask $M_{t}$ for each frame $f_{t}$. (2) A refined CLIP-based image encoder [gal2022image] $\mathcal{E}_{img}$ computes the cosine similarity between the feature embedding of the reference image and that of the masked subject in each frame. This targeted approach ensures the evaluation focuses strictly on identity fidelity by mitigating background interference. The final score is the average similarity over all $\mathbf{T}$ frames: $S_{\text{Nexus}}=\frac{1}{\mathbf{T}}\sum_{t=1}^{\mathbf{T}}\frac{\mathcal{E}_{\text{img}}(\mathbf{I}_{\text{ref}})\cdot\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})}{\|\mathcal{E}_{img}(\mathbf{I}_{\text{ref}})\|\|\mathcal{E}_{\text{img}}(f_{t}\odot M_{t})\|}$ where $\odot$ denotes element-wise multiplication.
@@
NaturalScore: This metric assesses the perceptual realism and physical plausibility of a generated video $\mathbf{V}$. We employ a state-of-the-art Multimodal Large Language Model (MLLM) to perform a deep semantic analysis. The model is presented with the video and a carefully designed prompt which instructs it to evaluate spatio-temporal consistency, identify violations of physical laws, and detect visual artifacts characteristic of generative models. The model’s direct output, a normalized score, serves as the final measure of the video’s naturalness.
@@
GmeScore: This metric evaluates the semantic alignment between a generated video $\mathbf{V}$ and its corresponding text prompt $\mathbf{C}$, proposed as an enhanced alternative to conventional CLIPScore. To overcome the limitations of prior methods in handling long-form text, GmeScore leverages the advanced text comprehension capabilities of the Qwen2.5-VL [Qwen2.5-VL], which provides a powerful text encoder $\mathcal{E}_{text}$ and a temporally-aware video encoder $\mathcal{E}_{video}$. Unlike methods that average per-frame similarities, GmeScore computes a global feature representation for the entire video, thereby capturing its holistic temporal dynamics. The final score is the cosine similarity: $S_{\text{Gme}}=\frac{\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\cdot\mathcal{E}_{\text{text}}(C)}{\|\mathcal{E}_{\text{video}}(\{f_{t}\}_{t=1}^{T})\|\|\mathcal{E}_{\text{text}}(C)\|}$ This formulation provides a more accurate and comprehensive assessment of text-to-video relevance, especially for complex narratives.
Appendix D Additional Results and Analysis